### v1

@@ 3.3. Well-posedness of the Bochner Laplacian in $\mathbf{H}^{1}_{t}(\Gamma)$
where $\mathbf{f}$ is a given everywhere tangential force and $\mathbf{P}:=\mathbf{I}-\boldsymbol{\nu}\otimes\boldsymbol{\nu}$ is the projection operator onto the tangent plane to $\Gamma$. We stress that the well-posedness and regularity results for the Laplace–Beltrami problem from the previous section do not immediately translate to (VL) due to the fact that $\Delta_{B}=\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}$ does not coincide with the component-wise application of the Laplace–Beltrami operator. In this section we establish the well-posedness of the $\mathrm{L}^{2}$-based weak formulation for (VL): find $\mathbf{u}\in\mathbf{H}^{1}_{t}(\Gamma)$ such that
$\int_{\Gamma}\nabla_{\Gamma}\mathbf{u}:\nabla_{\Gamma}\mathbf{v}=\langle\mathbf{f},\mathbf{v}\rangle,\quad\forall\,\mathbf{v}\in\mathbf{H}^{1}_{t}(\Gamma).$ (1.3)
We start by showing that replacing $\nabla_{M}$ by the (in principle) weaker $\nabla_{\Gamma}=\mathbf{P}\nabla_{M}$ yields an equivalent norm in $\mathbf{W}^{1,p}_{t}(\Gamma)$ for $1<p<\infty$ (and in particular, in $\mathbf{H}^{1}_{t}(\Gamma)=\mathbf{W}^{1,2}_{t}(\Gamma)$).
@@
Let $1<p<\infty$. There exists $C>0$ depending on $\Gamma$ and $p$ such that
$\|\nabla_{\Gamma}\mathbf{v}\|_{0,p;\Gamma}+\|\mathbf{v}\|_{0,p;\Gamma}\geq C\|\mathbf{v}\|_{1,p;\Gamma},\qquad\forall\mathbf{v}\in\mathbf{W}^{1,p}_{t}(\Gamma).$ (3.24)
@@
The proof is a simplified version of [40, Lemma 4.1]. From [40, eq. (2.17)] and 2.11 we know that
$\mathbf{B}\mathbf{v}=-(\nabla_{M}^{T}\mathbf{v})\boldsymbol{\nu},\qquad\forall\mathbf{v}\in\mathbf{W}^{1,p}_{t}(\Gamma).$ (3.25)
$\nabla_{M}\mathbf{v}=\nabla_{\Gamma}\mathbf{v}-\boldsymbol{\nu}\mathbf{v}^{T}\mathbf{B}.$ (3.26)
@@
By assuming further regularity on the manifold $\Gamma$ we will show that the lower order term $\|\mathbf{v}\|_{0,p;\Gamma}$ can actually be removed in the left-hand side of (3.24). This hinges on the injectivity of the operator $\nabla_{\Gamma}$ on $\mathbf{W}^{1,p}_{t}(\Gamma)$, which has been previously stated in [37, Lemma 2.1] for $p=2$ and a $C^{\infty}$ surface embedded in $\mathrm{R}^{3}$ (i.e. $d=2$).
The proof in [37, Lemma 2.1] relies on subtle differential geometry manipulations to first prove the injectivity of $\nabla_{\Gamma}$ on $C^{\infty}_{t}(\Gamma,\mathrm{R}^{3})$; then the proof asserts, without further justification, that said injectivity is transferred via density to $\mathbf{H}^{1}_{t}(\Gamma)$. This, however, is not straightforward and it does not follow from abstract functional analysis; there are non-injective linear continuous operators $T:X\rightarrow Y$ between two Hilbert spaces such that the restriction of $T$ to a dense subset of $X$ is injective (e.g. [66]). In the following lemma, we fill in this theoretical gap (see Step 1 within the proof below) and extend the result for less-regular manifolds of dimension $d\geq 2$ and $1<p<\infty$.
@@
Let $1<p<\infty$ and assume $\Gamma$ is $C^{3}$. Then, $\nabla_{\Gamma}$ is injective on $\mathbf{W}^{1,p}_{t}(\Gamma)$.
@@
Step 1: We start by showing that if $\mathbf{v}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ is such that $\nabla_{\Gamma}\mathbf{v}=\textbf{0}$ a.e. on $\Gamma$, then $\mathbf{v}\in\mathbf{C}^{2}_{t}(\Gamma)$.
Let $\mathbf{v}$ as above. Then, by (3.26) we know that
@@
Notice that the $C^{3}$-regularity of $\Gamma$ implies that $\boldsymbol{\nu}$ and $\mathbf{B}$ are respectively $C^{2}$ and $C^{1}$ functions. Thus, it is sufficient to show that $\mathbf{v}$ is merely continuous, as (3.27) would immediately imply the $C^{2}$-regularity of $\mathbf{v}$. The following Claim will prove useful:
Claim 3.7.
Let $n\in\mathrm{N}\cup\{0\}$ such that $pn<d$ and assume that $\mathbf{v}\in\mathbf{W}^{1,\frac{pd}{d-pn}}(\Gamma)$. Then, the following statements hold true:
  1. (i)
    If $p(n+1)<d$, then $\mathbf{v}\in\mathbf{W}^{1,\frac{pd}{d-p(n+1)}}(\Gamma)$.
  2. (ii)
    If $p(n+1)\geq d$, then $\mathbf{v}\in C(\Gamma)$.
Proof of 3.7. For (missing) i, since $\mathbf{v}\in\mathbf{W}^{1,\frac{pd}{d-pn}}(\Gamma)$ and $\frac{pd}{d-pn}<d$ (because $p(n+1)<d$) we have by 2.6 (Gagliardo–Nirenberg–Sobolev inequality) that $\mathbf{v}\in\mathbf{L}^{\frac{pd}{d-p(n+1)}}(\Gamma)$. This last fact combined with (3.27) yields $\mathbf{v}\in\mathbf{W}^{1,\frac{pd}{d-p(n+1)}}(\Gamma)$.
Regarding (missing) ii, we start by assuming $p(n+1)>d>pn$. Then, since $\mathbf{v}\in\mathbf{W}^{1,\frac{pd}{d-pn}}(\Gamma)$ and $\frac{pd}{d-pn}>d$ (because $p(n+1)>d$) we have by 2.7 (Sobolev embedding in Hölder spaces) that $\mathbf{v}\in\mathbf{C}^{0,\alpha}(\Gamma)\subseteq\mathbf{C}(\Gamma)$, where $\alpha:=1-\frac{d}{\frac{pd}{d-pn}}=\frac{p(n+1)-d}{p}\in(0,1)$. On the other hand, if $p(n+1)=d$, then since $\mathbf{v}\in\mathbf{W}^{1,\frac{pd}{d-pn}}(\Gamma)$ we have by 2.6 that $\mathbf{v}\in\mathbf{L}^{q}(\Gamma)$ for all $q\in(\frac{pd}{d-pn},\infty)=(d,\infty)$. This fact, combined with (3.27), yields $\mathbf{v}\in\mathbf{W}^{1,q}(\Gamma)$ for all $q\in(d,\infty)$. It is then enough to apply 2.7 with $q>d$ to obtain $\mathbf{v}\in\mathbf{C}^{0,\alpha}(\Gamma)\subseteq\mathbf{C}(\Gamma)$, where $\alpha:=1-\frac{d}{q}$. $\square$
Since $\mathbf{v}\in\mathbf{W}^{1,p}(\Gamma)=\mathbf{W}^{1,\frac{pd}{d-p\cdot 0}}(\Gamma)$, the proof of Step 1 then follows from iterating 3.7 starting with $n=0$.
@@
Step 2: Following Step 1 within [37, Lemma 2.1], we now aim to prove that: if $\mathbf{v}\in\mathbf{C}^{2}_{t}(\Gamma)$ is such that $\nabla_{\Gamma}\mathbf{v}=0$ on $\Gamma$, then $\mathbf{v}=\textbf{0}$ on a non-empty open set (relative to $\Gamma$).
Within the rest of the proof we write $\partial_{\mathbf{a}}\mathbf{v}:=(\nabla_{M}\mathbf{v})\mathbf{a}$, $D_{\mathbf{a}}\mathbf{v}:=(\nabla_{\Gamma}\mathbf{v})\mathbf{a}$ and $\mathbf{a},\mathbf{b}]:=\partial_{\mathbf{b}}\mathbf{a}-\partial_{\mathbf{a}}\mathbf{b}$.
The Riemannian Curvature tensor [21, §4.2, Definition 2.1] $R$ is defined by
$R(\mathbf{a},\mathbf{b},\mathbf{v}):=D_{\mathbf{a}}D_{\mathbf{b}}\mathbf{v}-D_{\mathbf{b}}D_{\mathbf{a}}\mathbf{v}-D_{[\mathbf{a},\mathbf{b}]}\mathbf{v}\in\mathbf{C}_{t}(\Gamma),$ (3.28)
@@
[…]
$R(\mathbf{a},\mathbf{b},\mathbf{v})=\mathbf{P}(\partial_{\mathbf{a}}\partial_{\mathbf{b}}\mathbf{v}-\partial_{\mathbf{b}}\partial_{\mathbf{a}}\mathbf{v}-\partial_{[\mathbf{a},\mathbf{b}]}\mathbf{v})+(\mathbf{v}^{T}\mathbf{B}\mathbf{b})\mathbf{B}\mathbf{a}-(\mathbf{v}^{T}\mathbf{B}\mathbf{a})\mathbf{B}\mathbf{b}.$ (3.29)
Furthermore, by taking into account the easily-verifiable identity
$(\partial_{\mathbf{a}}\partial_{\mathbf{b}}\mathbf{v})_{i}=\mathbf{b}^{T}\left\{(\nabla_{M}^{T}\nabla_{M}v_{i})\mathbf{a}+(\nabla_{M}^{T}\mathbf{a})(\nabla_{M}v_{i})\right\},$
and its analogue for $(\partial_{\mathbf{b}}\partial_{\mathbf{a}}\mathbf{v})_{i}$, and noticing that $\mathbf{a}^{T}(\nabla_{M}^{T}\nabla_{M}v_{i})\mathbf{b}=\mathbf{b}^{T}(\nabla_{M}^{T}\nabla_{M}v_{i})\mathbf{a}$ because of (2.7) and the tangentiality of $\mathbf{a}$ and $\mathbf{b}$, it is easy to see that $\partial_{\mathbf{a}}\partial_{\mathbf{b}}\mathbf{v}-\partial_{\mathbf{b}}\partial_{\mathbf{a}}\mathbf{v}=\partial_{[\mathbf{a},\mathbf{b}]}\mathbf{v}$; hence (3.29) reduces to
$R(\mathbf{a},\mathbf{b},\mathbf{v})=(\mathbf{v}^{T}\mathbf{B}\mathbf{b})\mathbf{B}\mathbf{a}-(\mathbf{v}^{T}\mathbf{B}\mathbf{a})\mathbf{B}\mathbf{b}.$ (3.30)
Fix $\mathbf{x}\in\Gamma$. Let $\{\mathbf{t}_{i}\}_{i=1}^{d}$ be a $C^{2}$ orthonormal basis of the tangent plane near $\mathbf{x}$; such basis can be obtained by applying the Gram-Schmidt method to the columns of $(\nabla\widetilde{\boldsymbol{\chi}})\circ\widetilde{\boldsymbol{\chi}}^{-1}$, where $\widetilde{\boldsymbol{\chi}}$ is a local parametrization of $\Gamma$ near $\mathbf{x}$. Denote by $\mathbf{T}\in\mathrm{R}^{(d+1)\times d}$ the matrix whose $i$-th column is the vector $\mathbf{t}_{i}$. Notice that we cannot select $\{\mathbf{t}_{i}\}_{i=1}^{d}$ to be the eigenframe of $\mathbf{B}$, as such eigenframe may be discontinuous at $\mathbf{x}$ if not all of the principal curvatures are simple [45]. From our purposes, we circumvent this technicality as follows: Recall that the action of $\mathbf{B}$ on the tangent plane of $\Gamma$ can be represented by the $d\times d$ tensor $\mathbf{A}:=\mathbf{T}^{T}\mathbf{B}\mathbf{T}$. Let $\{\mathbf{p}_{i}\}_{i=1}^{d}\subseteq\mathrm{R}^{d}$ be the eigenframe of $\widetilde{\mathbf{A}}$ at $\mathbf{y}:=\widetilde{\boldsymbol{\chi}}^{-1}(\mathbf{x})$; i.e. $\widetilde{\mathbf{A}}\mathbf{p}_{i}\big{\rvert}_{\mathbf{y}}=\widetilde{\kappa}_{i}\mathbf{p}_{i}\big{\rvert}_{\mathbf{y}}$, where $\{\kappa_{i}\}_{i=1}^{d}$ are the principal curvatures of $\Gamma$. Moreover, since $\mathbf{T}\mathbf{T}^{T}=\mathbf{I}-\boldsymbol{\nu}\otimes\boldsymbol{\nu}$ is the projection onto the tangent plane of $\Gamma$, we have that $\widetilde{\mathbf{B}}(\widetilde{\mathbf{T}}\mathbf{p}_{i})\big{\rvert}_{\mathbf{y}}=\widetilde{\mathbf{T}}(\widetilde{\mathbf{T}}^{T}\widetilde{\mathbf{B}}\widetilde{\mathbf{T}})\mathbf{p}_{i}\big{\rvert}_{\mathbf{y}}=\widetilde{\kappa}_{i}(\widetilde{\mathbf{T}}\mathbf{p}_{i})\big{\rvert}_{\mathbf{y}}$. Hence, $\{\mathbf{r}_{i}:=\mathbf{T}\mathbf{p}_{i}\}_{i=1}^{d}$ is also a $C^{2}$ orthonormal basis of the tangent plane near $\mathbf{x}$ and $\mathbf{B}\mathbf{r}_{i}\big{\rvert}_{\mathbf{x}}=\kappa_{i}\mathbf{r}_{i}\big{\rvert}_{\mathbf{x}}$ for each $i\in\{1,\dotsc,d\}$.
Invoking (3.30), elementary calculations then show that
$\left[R(\mathbf{r}_{i},\mathbf{r}_{j},\mathbf{r}_{k})\cdot\mathbf{r}_{i}\right]\big{\rvert}_{\mathbf{x}}=\delta_{kj}\kappa_{i}(\mathbf{x})\kappa_{j}(\mathbf{x}),\qquad\forall i,j,k=1,\dotsc,d,\quad i\neq j.$ (3.31)
Now, let $\mathbf{v}\in\mathbf{C}^{2}_{t}(\Gamma)$ be such that $\nabla_{\Gamma}\mathbf{v}=\textbf{0}$ on $\Gamma$, whence $R(\mathbf{r}_{i},\mathbf{r}_{j},\mathbf{v})=0$ according to (3.28). We express $\mathbf{v}$ locally near $\mathbf{x}$ by $\mathbf{v}=\sum_{k=1}^{d}\alpha_{k}\mathbf{r}_{k}$. Then, using (3.31) we have for each $i,j=1,\dotsc,d$ with $i\neq j$,
$0=\left[R(\mathbf{r}_{i},\mathbf{r}_{j},\mathbf{v})\cdot\mathbf{r}_{i}\right]\big{\rvert}_{\mathbf{x}}=\sum_{k=1}^{d}\alpha_{k}(\mathbf{x})\left[R(\mathbf{r}_{i},\mathbf{r}_{j},\mathbf{r}_{k})\cdot\mathbf{r}_{i}\right]\big{\rvert}_{\mathbf{x}}=\sum_{k=1}^{d}\alpha_{k}(\mathbf{x})\delta_{kj}\kappa_{i}(\mathbf{x})\kappa_{j}(\mathbf{x})=\alpha_{j}(\mathbf{x})\kappa_{i}(\mathbf{x})\kappa_{j}(\mathbf{x}).$ (3.32)
@@
Since $\Gamma$ is an $d$-dimensional compact manifold without boundary embedded in $\mathrm{R}^{d+1}$, we have by [76, Theorem 4, p.88] (still valid for $C^{2}$ surfaces [68]) and the continuity of the Gauss curvature $K:=\prod_{i=1}^{d}\kappa_{i}$ on $\Gamma$, that the set $\{\mathbf{z}\in\Gamma:K(\mathbf{z})\neq 0\}=\{\mathbf{z}\in\Gamma:\kappa_{i}(\mathbf{z})\neq 0,\quad\forall\,1,\dotsc,d\}$ is a non-empty open set (relative to $\Gamma$). For each point $\mathbf{x}$ in that set, identity (3.32) yields $\alpha_{j}(\mathbf{x})=0$ for each $j=1,\dotsc,d$, whence $\mathbf{v}(\mathbf{x})=\sum_{k=1}^{d}\alpha_{k}(\mathbf{x})\mathbf{r}_{k}(\mathbf{x})=\textbf{0}$.
@@
Step 3: Let $\mathbf{v}\in\mathbf{C}^{2}_{t}(\Gamma)$ be such that $\nabla_{\Gamma}\mathbf{v}=0$ on $\Gamma$ and there is $\mathbf{x}\in\Gamma$ such that $\mathbf{v}(\mathbf{x})=0$. We now aim to prove that $\mathbf{v}(\overline{\mathbf{x}})=0$ for each $\overline{\mathbf{x}}\in\Gamma$.
@@
The proof will follow almost verbatim to Step 2 within [37, Lemma 2.1]. Given that $\Gamma$ is not necessarily smooth, but only $C^{3}$, the existence of a geodesic on $\Gamma$ connecting $\mathbf{x}$ and $\overline{\mathbf{x}}$ does not follow from the classical Hopf–Rinow theorem from Riemannian geometry [21, Chapter 7, Theorem 2.8]. However, by [14, Chapter I, Proposition 3.18] (which holds for $C^{1}$ manifolds), we deduce that $(\Gamma,d)$ is a length space where $d$ is the metric on $\Gamma$ defined as follows: for each $\mathbf{z},\mathbf{w}\in\Gamma$, let $d(\mathbf{z},\mathbf{w})$ be the infimum of $\int_{0}^{1}|\dot{\boldsymbol{\gamma}}(t)|\,{\mathrm{d}}t$ over all piecewise-$C^{1}$ paths $\boldsymbol{\gamma}:[0,1]\rightarrow\Gamma$ such that $\boldsymbol{\gamma}(0)=\mathbf{z}$ and $\boldsymbol{\gamma}(1)=\mathbf{w}$. Thus, by the Hopf–Rinow theorem in length spaces [14, Chapter I, Proposition 3.7], [15, Theorem 2.5.28], we deduce that $(\Gamma,d)$ is geodesically complete; i.e. the infimum defining $d$ is realizable by a (in principle) piecewise-$C^{1}$ path $\boldsymbol{\gamma}$ called geodesic. Upon reparametrizing $\boldsymbol{\gamma}$ by its arc-length (so it has constant speed), we deduce from [77, Proposition 1.1] and [77, Corollary 2.7] that $\boldsymbol{\gamma}$ is $C^{1}$. The rest of the proof follows as in [37, Lemma 2.1], which we describe below for the sake of completeness.

### v2

@@ 3.1. $\mathrm{L}^{p}$-based Sobolev regularity
where $\mathbf{f}$ is a given everywhere tangential force and $\mathbf{P}:=\mathbf{I}-\boldsymbol{\nu}\otimes\boldsymbol{\nu}$ is the projection operator onto the tangent plane to $\Gamma$. We stress that the well-posedness and regularity results for the Laplace–Beltrami problem from the previous section do not immediately translate to (VL) due to the fact that $\Delta_{B}=\mathbf{P}\operatorname{div}_{\Gamma}\nabla_{\Gamma}=\mathbf{P}\operatorname{div}_{M}(\mathbf{P}\nabla_{M})$ does not coincide with the component-wise application of the Laplace–Beltrami operator $\operatorname{div}_{M}\nabla_{M}$. In this section we are concerned with the well-posedness of the $\mathrm{L}^{p}$-based ($1<p<\infty$) weak formulation for (VL): find $\mathbf{u}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ such that
We start by showing that replacing $\nabla_{M}$ by the (in principle) weaker $\nabla_{\Gamma}=\mathbf{P}\nabla_{M}$ yields an equivalent norm in $\mathbf{W}^{1,p}_{t}(\Gamma)$ for $p\in(1,\infty)$ (and in particular, in $\mathbf{H}^{1}_{t}(\Gamma)=\mathbf{W}^{1,2}_{t}(\Gamma)$).
@@
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{2}$. Then,
$\|\nabla_{\Gamma}\mathbf{v}\|_{0,p;\Gamma}+\|\mathbf{v}\|_{0,p;\Gamma}\geq C\|\mathbf{v}\|_{1,p;\Gamma},\qquad\forall\mathbf{v}\in\mathbf{W}^{1,p}_{t}(\Gamma),$ (3.3)
where the constant $C$ depends only on $p$ and linearly on $\|\mathbf{B}\|_{0;\infty;\Gamma}$.
@@
The proof is a simplified version of [36, Lemma 4.1]. From (2.7) and 2.7 (density) we know that
$\nabla_{M}\mathbf{v}=\nabla_{\Gamma}\mathbf{v}-\boldsymbol{\nu}\mathbf{v}^{T}\mathbf{B},\qquad\forall\mathbf{v}\in\mathbf{W}^{1,p}_{t}(\Gamma).$
@@
We will show that the lower order term $\|\mathbf{v}\|_{0,p;\Gamma}$ can actually be removed in the left-hand side of (3.3). This hinges on the injectivity of the operator $\nabla_{\Gamma}$ on $\mathbf{W}^{1,p}_{t}(\Gamma)$, which has been previously stated in [34, Lemma 2.1] for $p=2$ and a $C^{\infty}$ surface embedded in $\mathrm{R}^{3}$ (i.e. $d=2$).
The proof in [34, Lemma 2.1] relies on subtle differential geometry manipulations to first prove the injectivity of $\nabla_{\Gamma}$ on $\mathbf{C}^{\infty}_{t}(\Gamma)$; then the proof asserts, without further justification, that said injectivity is transferred via density to $\mathbf{H}^{1}_{t}(\Gamma)$. This, however, is not straightforward and it does not follow from abstract functional analysis; there are non-injective linear continuous operators $T:X\rightarrow Y$ between two Hilbert spaces such that the restriction of $T$ to a dense subset of $X$ is injective (e.g. [57]). In the following lemma, we fill in this theoretical gap (see Step 1 within the proof below) and extend the result for less-regular manifolds of dimension $d\geq 2$ and $p\in(1,\infty)$. Moreover, we note that the “weak” Riemann curvature tensor (cf. (C.3)) is instrumental to the proof.
@@
Let $p\in(1,\infty)$ and assume $\Gamma$ is of class $C^{2}$. Then, $\nabla_{\Gamma}$ is injective on $\mathbf{W}^{1,p}_{t}(\Gamma)$.
@@
Step 1: We start by showing that if $\mathbf{v}\in\mathbf{W}^{1,p}_{t}(\Gamma)$ is such that $\nabla_{\Gamma}\mathbf{v}=\textbf{0}$ a.e. on $\Gamma$, then $\mathbf{v}\in\mathbf{C}^{1}_{t}(\Gamma)$. Let $\mathbf{v}$ as above. Then, by (2.7) we know that
@@
If $\mathbf{v}\in\mathbf{W}^{1,p}_{t}(\Gamma)$, then by 2.8 (Gagliardo–Nirenberg–Sobolev inequality), it transpires that $\mathbf{v}\in\mathbf{L}^{p_{0}}_{t}(\Gamma)$ for $p_{0}=\frac{pd}{d-p(n+1)}$. Notice that the $C^{2}$-regularity of $\Gamma$ implies that $\boldsymbol{\nu}$ and $\mathbf{B}$ are respectively $C^{1}$ and $C^{0}$ functions and so from (3.4) we deduce that $\mathbf{v}\in\mathbf{L}^{p_{0}}(\Gamma)$ implies $\nabla_{M}\mathbf{v}\in\mathbb{L}^{p_{0}}(\Gamma)$. Iterating this process leads to sufficiently high integrability of $\mathbf{v}$ so as to use 2.9 (Sobolev embedding in Hölder spaces) to obtain $\mathbf{v}\in\mathbf{C}^{0}_{t}(\Gamma)$, and so, again, by (3.4), $\mathbf{v}\in\mathbf{C}^{1}_{t}(\Gamma)$. This completes the proof of Step 1.
@@
Step 2: We claim that: if $\mathbf{v}\in\mathbf{C}^{1}_{t}(\Gamma)$ is such that $\nabla_{\Gamma}\mathbf{v}=\textbf{0}$ on $\Gamma$, then $\mathbf{v}=\textbf{0}$ on a non-empty open set (relative to $\Gamma$).
Fix $\mathbf{x}\in\Gamma$. Let $\{\mathbf{t}_{i}\}_{i=1}^{d}$ be a $C^{1}$ orthonormal basis of the tangent plane near $\mathbf{x}$; such basis can be obtained by applying the Gram–Schmidt method to the columns of $(\nabla\widetilde{\boldsymbol{\chi}})\circ\widetilde{\boldsymbol{\chi}}^{-1}$, where $\widetilde{\boldsymbol{\chi}}$ is a local parametrization of $\Gamma$ near $\mathbf{x}$. Denote by $\mathbf{T}\in\mathrm{R}^{(d+1)\times d}$ the matrix field whose $i$-th column is $\mathbf{t}_{i}$. Notice that we cannot select $\{\mathbf{t}_{i}\}_{i=1}^{d}$ to be the eigenframe of $\mathbf{B}$, as such eigenframe may be discontinuous at $\mathbf{x}$ if not all of the principal curvatures are simple [41]. From our purposes, we circumvent this technicality as follows: recall that the action of $\mathbf{B}$ on the tangent plane of $\Gamma$ can be represented by the $d\times d$ tensor $\mathbf{A}:=\mathbf{T}^{T}\mathbf{B}\mathbf{T}$. Let $\{\mathbf{p}_{i}\}_{i=1}^{d}\subseteq\mathrm{R}^{d}$ be the eigenframe of $\widetilde{\mathbf{A}}$ at $\mathbf{y}:=\widetilde{\boldsymbol{\chi}}^{-1}(\mathbf{x})$; i.e. $\widetilde{\mathbf{A}}\mathbf{p}_{i}\big\rvert_{\mathbf{y}}=\widetilde{\kappa}_{i}\mathbf{p}_{i}\big\rvert_{\mathbf{y}}$, where $\{\kappa_{i}\}_{i=1}^{d}$ are the principal curvatures of $\Gamma$. Moreover, since $\mathbf{T}\mathbf{T}^{T}=\mathbf{I}-\boldsymbol{\nu}\otimes\boldsymbol{\nu}$ is the projection onto the tangent plane of $\Gamma$, we have that $\widetilde{\mathbf{B}}(\widetilde{\mathbf{T}}\mathbf{p}_{i})\big\rvert_{\mathbf{y}}=\widetilde{\mathbf{T}}(\widetilde{\mathbf{T}}^{T}\widetilde{\mathbf{B}}\widetilde{\mathbf{T}})\mathbf{p}_{i}\big\rvert_{\mathbf{y}}=\widetilde{\kappa}_{i}(\widetilde{\mathbf{T}}\mathbf{p}_{i})\big\rvert_{\mathbf{y}}$. Hence, $\{\mathbf{r}_{i}:=\mathbf{T}\mathbf{p}_{i}\}_{i=1}^{d}$ is also a $C^{1}$ orthonormal basis of the tangent plane near $\mathbf{x}$ and $\mathbf{B}\mathbf{r}_{i}\big\rvert_{\mathbf{x}}=\kappa_{i}\mathbf{r}_{i}\big\rvert_{\mathbf{x}}$ for each $i\in\{1,\dotsc,d\}$.
Now, let $\mathbf{v}\in\mathbf{C}^{1}_{t}(\Gamma)$ be such that $\nabla_{\Gamma}\mathbf{v}=\textbf{0}$ on $\Gamma$. From (C.3) and (C.6) (Riemann Curvature Tensor) and a localization argument, we deduce that for each $\mathbf{a},\mathbf{b}\in\mathbf{C}^{1}_{t}(\Gamma)$ it holds that
@@
By expressing $\mathbf{v}$ locally near $\mathbf{x}$ by $\mathbf{v}=\sum_{k=1}^{d}\alpha_{k}\mathbf{r}_{k}$, it follows from (3.5) that for each $i,j=1,\dotsc,d$:
$0=\left[(\mathbf{v}^{T}\mathbf{B}\mathbf{r}_{j})\mathbf{B}\mathbf{r}_{i}-(\mathbf{v}^{T}\mathbf{B}\mathbf{r}_{i})\mathbf{B}\mathbf{r}_{j}\right]\big\rvert_{\mathbf{x}}=\left[\kappa_{i}\kappa_{j}\left((\mathbf{v}\cdot\mathbf{r}_{j})\mathbf{r}_{i}-(\mathbf{v}\cdot\mathbf{r}_{i})\mathbf{r}_{j}\right)\right]\big\rvert_{\mathbf{x}}=\left[\kappa_{i}\kappa_{j}\left(\alpha_{j}\mathbf{r}_{i}-\alpha_{i}\mathbf{r}_{j}\right)\right]\big\rvert_{\mathbf{x}};$
hence,
$\alpha_{j}(\mathbf{x})\kappa_{i}(\mathbf{x})\kappa_{j}(\mathbf{x})=0,\qquad\forall i,j=1,\dotsc,d.$ (3.6)
@@
Since $\Gamma$ is an $d$-dimensional compact manifold without boundary embedded in $\mathrm{R}^{d+1}$, we have by [64, Theorem 4, p.88] (still valid for $C^{2}$ manifolds [59]) and the continuity of the Gauss curvature $K:=\prod_{i=1}^{d}\kappa_{i}$ on $\Gamma$, that the set $\{\mathbf{z}\in\Gamma:K(\mathbf{z})\neq 0\}=\{\mathbf{z}\in\Gamma:\kappa_{i}(\mathbf{z})\neq 0,\quad\forall\,1,\dotsc,d\}$ is a non-empty open set (relative to $\Gamma$). For each point $\mathbf{x}$ in that set, identity (3.6) yields $\alpha_{j}(\mathbf{x})=0$ for each $j=1,\dotsc,d$, whence $\mathbf{v}(\mathbf{x})=\sum_{k=1}^{d}\alpha_{k}(\mathbf{x})\mathbf{r}_{k}(\mathbf{x})=\textbf{0}$.
@@
Step 3: Let $\mathbf{v}\in\mathbf{C}^{1}_{t}(\Gamma)$ be such that $\nabla_{\Gamma}\mathbf{v}=\textbf{0}$ on $\Gamma$ and there is $\mathbf{x}\in\Gamma$ such that $\mathbf{v}(\mathbf{x})=\textbf{0}$. We now aim to prove that $\mathbf{v}(\overline{\mathbf{x}})=\textbf{0}$ for each $\overline{\mathbf{x}}\in\Gamma$.
@@
Given that $\Gamma$ is not necessarily smooth, but only of class $C^{2}$ and connected, the existence of a geodesic on $\Gamma$ connecting $\mathbf{x}$ and $\overline{\mathbf{x}}$ does not follow from the classical Hopf–Rinow theorem from Riemannian geometry [20, Chapter 7, Theorem 2.8]. However, by [14, Chapter I, Proposition 3.18] (which holds for manifolds of class $C^{1}$), we deduce that $(\Gamma,d)$ is a length space where $d$ is the metric on $\Gamma$ defined as follows: for each $\mathbf{z},\mathbf{w}\in\Gamma$, let $d(\mathbf{z},\mathbf{w})$ be the infimum of $\int_{0}^{1}|\dot{\boldsymbol{\gamma}}(t)|\,{\mathrm{d}}t$ over all piecewise-$C^{1}$ paths $\boldsymbol{\gamma}:[0,1]\rightarrow\Gamma$ such that $\boldsymbol{\gamma}(0)=\mathbf{z}$ and $\boldsymbol{\gamma}(1)=\mathbf{w}$. Thus, by the Hopf–Rinow theorem in length spaces [14, Chapter I, Proposition 3.7], [15, Theorem 2.5.28], we deduce that $(\Gamma,d)$ is geodesically complete; i.e. the infimum defining $d$ is realizable by a (in principle) piecewise-$C^{1}$ path $\boldsymbol{\gamma}$ called geodesic. Upon reparametrizing $\boldsymbol{\gamma}$ by its arc-length (so it has constant speed), we deduce from [65, Proposition 1.1] and [65, Corollary 2.7] that $\boldsymbol{\gamma}$ is of class $C^{1}$.